OpenStudy (anonymous):

A square with a side of 18 is decreased to a side of 15. By how much is the diagonal decreased?

OpenStudy (anonymous):



OpenStudy (anonymous):

Find the diagonal of the big square: 18^2 + 18^2 = c^2 sqrt(18^2 + 18^2) = c Then find the diagonal of the small square: 15^2 + 15^2 = C^2 sqrt( 15^2 + 15^2) = C Then subtract to find the difference... try that.